# DualOptim: Enhancing Efficacy and Stability in Machine Unlearning with Dual Optimizers

### v1

@@ Abstract
Existing machine unlearning (MU) approaches exhibit significant sensitivity to hyperparameters, requiring meticulous tuning that limits practical deployment. In this work, we first empirically demonstrate the instability and suboptimal performance of existing popular MU methods when deployed in different scenarios. To address this issue, we propose Dual Optimizer (DualOptim), which incorporates adaptive learning rate and decoupled momentum factors. Empirical and theoretical evidence demonstrates that DualOptim contributes to effective and stable unlearning. Through extensive experiments, we show that DualOptim can significantly boost MU efficacy and stability across diverse tasks, including image classification, image generation, and large language models, making it a versatile approach to empower existing MU algorithms.
@@ 1 Introduction
Recent advancements in machine unlearning (MU) research have established it as a crucial technique for utilizing pretrained models while addressing various trustworthy challenges in various applications [1, 2]. These MU methods have two major categories: exact MU [3] and approximate MU [4]. Exact MU requires retraining a model on a dataset excluding forgetting samples from scratch, which is computationally prohibitive for large-scale models and datasets. Therefore, we study approximate MU methods in this work, with focuses on their efficacy and stability.
Most existing approximate MU methods [5, 6, 7, 8, 9] aim to maximize loss on forget samples while preserving performance on retain samples. While effective, these methods exhibit significant sensitivity to hyperparameter choices, with optimal configurations requiring meticulous tuning that varies substantially across datasets and forgetting scenarios. This dependency complicates their practical deployment. As illustrated in Figure 1, we empirically demonstrate that state-of-the-art MU approaches [6, 7, 8] exhibit suboptimal performance or instability. These limitations highlight the critical need for techniques to improve the effectiveness and stability of MU methods.
In this work, we first reveal the unstable performance of existing MU methods when they are applied in different scenarios. Considering the need to optimize two objectives on the forgetting and the retaining samples, we then propose Dual Optimizer (DualOptim) to address this challenge. Specifically, we employ an optimizer with adaptive learning rate, such as Adam [10, 11], to optimize the forgetting objective, while utilizing a separate optimizer for the retaining objective. Compared with existing methods, DualOptim decouples the momentum terms for the two objectives by using and updating them separately. In addition, DualOptim is a plug-and-play solution that can be easily integrated into existing MU algorithms.
@@
Figure 1: Unlearning process of MU baselines. SFRon [7], SalUn [6] and SCRUB [8] are adopted as the baselines. The metrics are those mentioned in Sec. 3.1. All results are obtained from unlearning 10% random subset of CIFAR-10 on ResNet-18. The solid lines and shadows denote the mean and standard deviation across 5 trials with different random data. The hyperparameters of different methods are selected based on minimizing the averaging gap between retraining and them across 5 trials. The red dashed lines denote the final performance of retraining as a reference.
Our theoretical analysis validates that decoupling momentum terms in two separate optimizers can help decrease the variance of model parameters, indicating a more stable performance. Through comprehensive experiments, we validate the effectiveness of DualOptim in improving the performance of existing MU methods across various tasks, including image classification, image generation, and large language models. Our proposed technique is generic and capable of pushing the state-of-the-art performance on multiple tasks.
We summarize the contributions of this paper as follows:
@@ 2 Related Works
Machine Unlearning (MU). MU targets the need to remove specific data influences from pretrained models [12, 13, 14], while complying with privacy requirements tied to differential privacy [3, 15, 16, 17, 18]. Initially developed on linear classifiers with convex loss objective functions, exact MU [3, 15, 19, 20] allowed precise data removal under privacy budgets to counter privacy attacks. However, exact MU cannot be applied to deep neural networks due to their non-convex loss functions and prohibitive computational cost for full retraining. In this context, approximate MU [4, 21] was proposed to address this issue by fine-tuning the model to achieve the forgetting effect. Despite improved efficiency, approximate MU may cause catastrophic performance declines on retaining data [21, 22]. Recent methods incorporate fine-tuning [6, 8, 23], sparsity regularization [24], knowledge distillation [8, 25], saliency map [6, 7] and alternative updating [7] to better balance efficient forgetting and model utility. However, as illustrated in Figure 1, these methods generally exhibit suboptimal performance or high hyperparameter sensitivity, underscoring the necessity to develop a method to enhance efficacy and stability during unlearning.
MU for Multiple Tasks. Besides classification, MU has broad applications for multiple tasks. For example, recent text-conditional image generation models have showcased their ability to produce images that closely align with textual descriptions [26, 27, 28, 29]. However, significant security and privacy concerns have been raised [1, 30, 31], necessitating the application of MU methods to these models. While early works [30, 32, 33, 34] focuses on concept deletion in diffusion models, recent studies [6, 7, 35] improve their performance and propose methods applicable to more general image generators. Another notable example is large language models (LLMs), which have demonstrated remarkable capabilities [36, 37] but also face privacy and copyright issues like retaining unauthorized content [38, 39, 40, 41]. To address these concerns, MU methods have been employed to fine-tune LLMs [2, 9, 42] to effectively and efficiently achieve data forgetting. However, all these methods struggle to achieve a balance between model utility and forget effectiveness. In addition, intensive hyperparameter tuning is expected for optimal performance, bringing challenges for practitioners.
@@ 3.1 Preliminary
Let $\mathcal{D}=\{{\bm{z}}_{i}\}_{i=1}^{N}$ denote the training set for pretraining. The subset of the training set we aim to forget during unlearning is known as the forget set $\mathcal{D}_{f}\subset\mathcal{D}$, and its complement, $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$ is the retain set. In the context of MU, we denote the parameters of pretrained model as $\theta_{o}\in{\mathbb{R}}^{d}$, which is trained on $\mathcal{D}$. Consistent with previous studies [6, 7, 21, 24], Retraining is considered the gold standard for MU, where the model parameters $\theta_{r}\in\mathbb{R}^{d}$ are trained from scratch on $\mathcal{D}_{r}$. However, retraining is computationally intensive or even infeasible, especially for large models and datasets. This poses a significant challenge in the practical applications of MU.
@@
We focus on approximate MU in this work, its primary objective is to obtain an unlearned model, referred to as $\theta_{u}\in{\mathbb{R}}^{d}$, from $\theta_{o}$ on $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$ so that it can serve as an accurate and computationally efficient alternative to the retrained model $\theta_{r}$. Mathematically, MU aims to solve the optimization problem: $\min_{\theta}\mathcal{L}_{f}(\mathcal{D}_{f},\theta)+\mathcal{L}_{r}(\mathcal{D}_{r},\theta)$, where $\mathcal{L}_{f}$ and $\mathcal{L}_{r}$ denote forget and retain loss objective functions, respectively. In practice, ${\mathcal{L}}_{r}$ is usually the same as the loss objective function in pre-training and ${\mathcal{L}}_{f}$ is the opposite, so MU aims to improve the performance on ${\mathcal{D}}_{r}$ while degrading the performance on ${\mathcal{D}}_{f}$. To avoid confusion, we call ${\mathcal{L}}_{f}({\mathcal{D}}_{f},\theta)$, ${\mathcal{L}}_{r}({\mathcal{D}}_{r},\theta)$ that we aim to minimize forget loss and retain loss, respectively. In addition, many MU algorithms [6, 7, 8] update $\theta$ by alternately optimizing $\mathcal{L}_{f}({\mathcal{D}}_{f},\theta)$ and $\mathcal{L}_{r}({\mathcal{D}}_{r},\theta)$ for better performance. The generic pipeline of MU is presented in Algorithm 1. For notation simplicity, we use ${\bm{g}}_{f}:=\nabla_{\theta}{\mathcal{L}}_{f}({\mathcal{D}}_{f},\theta)$ and ${\bm{g}}_{r}:=\nabla_{\theta}{\mathcal{L}}_{r}({\mathcal{D}}_{r},\theta)$ to denote the gradients of the forget loss and the retain loss, respectively. In mini-batch updates, we use $\widehat{{\bm{g}}}_{f}$ and $\widehat{{\bm{g}}}_{r}$ to denote the corresponding stochastic gradients.
@@
The numerical analysis in this section is based on classification tasks. We adopt the same evaluation schemes as [7]. That is, we use the accuracy on the forget set (FA), the retain set (RA), test set (TA) and the success rate of membership inference attack (MIA) [43]  [^2] We use the same implementation as [7] and adopt the entropy of the output probabilities as the metric in MIA. on the forgetting set to indicate model utility and the forgetting effectiveness. A competitive unlearning method should have stable performance and small gaps with retraining in all these four evaluation schemes.
@@ 3.2 Adaptive Learning Rate Enables Stable Forgetting
Despite the effectiveness of existing MU methods, achieving satisfactory unlearning performance often relies on carefully selecting hyperparameters. The optimal hyperparameters can vary significantly across different forget sets, complicating the practical applications of MU algorithms. As shown in Figure 2 (a)-(d), we use the same hyperparameters for SFRon [7], the best-performed baseline in our evaluation, with the default SGD optimizer for 5 trials of different randomly sampled forget sets. We defer more results of other existing methods to Appendix C.6. Despite being intensively tuned, the unified hyperparameters exhibit quite unstable performance across different forget sets during unlearning. This underscores that a hyperparameter configuration effective for certain forget sets may not be suitable for others. When dealing with a new forget set, even if it is sampled in the same way, the need for precise hyperparameter tuning can pose challenges for practitioners.
@@
Figure 2: Unlearning process with different ablations of the proposed method. All results are obtained from unlearning 10% random subset of CIFAR-10 by SFRon [7] on ResNet-18. (a)-(d) The metrics are those mentioned in Sec. 3.1. The red dashed lines denote the final performance of retraining as a reference. The solid lines and shadows denote the mean and standard deviation across 5 trials with different random forget sets.
Figure 3 (a) demonstrates the magnitude of the gradients on the forget set and the retain set when using SGD optimizer. We can clearly see that the gradients on the forget set have significantly larger magnitudes with substantial variations compared with the ones on the retain set. In this context, it is crucial to adaptively adjust the learning rate to handle different gradient magnitudes. Therefore, we use optimizers with preconditioners, such as Adam [10, 11] to evade tricky hyperparameter tuning for the learning rate. Specifically, Adam employs adaptive learning rates based on the historic gradient magnitudes, making it suitable to handle large gradient magnitude variations during unlearning.
@@
The observations in Figure 2 (a)-(d) and Figure 3 (b) suggest that Adam provides more stable performance and induces smaller gradient norm across different forget sets. However, despite enhanced stability, the noticeable performance gap with retraining, especially in the metric of membership inference attack (MIA), indicates that Adam only achieves suboptimal unlearning efficacy. We need a mechanism to improve the performance while ensuring the stability of the method.
@@ 3.3 Decoupled Momentum for Enhanced Stability in Machine Unlearning
Inspired by the disparities in the magnitudes of $\widehat{{\bm{g}}}_{f}$ and $\widehat{{\bm{g}}}_{r}$, along with their low correlation observed in Figure 3 and 4, we introduce decoupled momentum, which employs two separate momentum terms dedicated to the forget loss ${\mathcal{L}}_{f}$ and the retain loss ${\mathcal{L}}_{r}$, respectively. This approach ensures that the momentum update for the forget loss remains unaffected by the gradients of the retaining loss, and vice versa. By decoupling the momentum terms in this manner, we can more effectively optimize the distinct processes of forgetting and retaining, leading to enhanced stability and performance. As illustrated in Figure 2 (a)-(d) (represented by the green lines) and Figure 3 (c), the combination of adaptive learning rate with decoupled momentum not only stabilizes the unlearning process but also leads to improved unlearning efficacy. Note that as presented in Appendix C.6, although other methods like SalUn [6] and SCRUB [8] perform more stable than SFRon, they only achieve suboptimal unlearning performance. Nevertheless, our method can also significantly improve their performance.
@@
(Correlation Bounds) The correlation between the stochastic gradients from the same function in different time steps is bounded while the correlation between stochastic gradients from different functions can be ignored. That is to say, $\exists\tau\in[0,1]$ such that:
@@
The assumption $\forall t_{1},t_{2}$, $\rho(\widehat{{\bm{g}}}_{f,t_{1}},\widehat{{\bm{g}}}_{r,t_{2}})\simeq 0$ is motivated by the observations in Figure 4 of Appendix C.1 and for sake of notation simplicity. Our analyses can be easily extended to the cases if we use a small constant to bound this correlation.
@@
$\mathrm{(Shared~{}Momentum)}\begin{cases}{\bm{m}}_{f,t}^{S}&=\alpha{\bm{m}}_{r,t-1}^{S}+\widehat{{\bm{g}}}^{S}_{f,t},\quad\theta^{S}_{f,t}=\theta^{S}_{r,t-1}-\eta{\bm{m}}_{f,t}^{S}\\ {\bm{m}}_{r,t}^{S}&=\alpha{\bm{m}}_{f,t}^{S}+\widehat{{\bm{g}}}^{S}_{r,t},\quad\quad\theta^{S}_{r,t}=\theta^{S}_{f,t}-\eta{\bm{m}}_{r,t}^{S}\end{cases}$ (5)
$\mathrm{(Decoupled~{}Momentum)}\begin{cases}{\bm{m}}_{f,t}^{D}&=\alpha{\bm{m}}_{f,t-1}^{D}+\widehat{{\bm{g}}}^{D}_{f,t},\quad\theta^{D}_{f,t}=\theta^{D}_{r,t-1}-\eta{\bm{m}}_{f,t}^{D}\\ {\bm{m}}_{r,t}^{D}&=\alpha{\bm{m}}_{r,t-1}^{D}+\widehat{{\bm{g}}}^{D}_{r,t},\quad~{}\theta^{D}_{r,t}=\theta^{D}_{f,t}-\eta{\bm{m}}_{r,t}^{D}\end{cases}$
@@
The proof is deferred to Appendix A.2. Theorem 3.5 formalizes that decoupled momentum induces smaller worst-case parameter variance compared to shared momentum, demonstrating that decoupled momentum theoretically enhances the stability of unlearning.

### v2

@@ Abstract
Existing machine unlearning (MU) approaches exhibit significant sensitivity to hyperparameters, requiring meticulous tuning that limits practical deployment. In this work, we first empirically demonstrate the instability and suboptimal performance of existing popular MU methods when deployed in different scenarios. To address this issue, we propose Dual Optimizer (DualOptim), which incorporates adaptive learning rate and decoupled momentum factors. Empirical and theoretical evidence demonstrates that DualOptim contributes to effective and stable unlearning. Through extensive experiments, we show that DualOptim can significantly boost MU efficacy and stability across diverse tasks, including image classification, image generation, and large language models, making it a versatile approach to empower existing MU algorithms. Codes are available at https://github.com/CityU-MLO/DualOptim.
@@ 1 Introduction
Recent advancements in machine unlearning (MU) research have established it as a crucial technique for utilizing pretrained models while addressing various trustworthy challenges in various applications randored ; mainitofu . These MU methods have two major categories: exact MU (guo2020certified, ) and approximate MU (izzo2021approximate, ). Exact MU requires retraining a model on a dataset excluding forgetting samples from scratch, which is computationally prohibitive for large-scale models and datasets. Therefore, we study approximate MU methods in this work, with focuses on their efficacy and stability.
Most existing approximate MU methods (graves2021amnesiac, ; fan2024salun, ; huang2025unified, ; kurmanji2023towards, ; yuan2024closer, ) aim to maximize loss on forget samples while preserving performance on retain samples. While effective, these methods exhibit significant sensitivity to hyperparameter choices, with optimal configurations requiring meticulous tuning that varies substantially across datasets and forgetting scenarios. This dependency complicates their practical deployment. As illustrated in Figure 1, we empirically demonstrate that state-of-the-art MU approaches (fan2024salun, ; huang2025unified, ; kurmanji2023towards, ) exhibit suboptimal performance or instability. These limitations highlight the critical need for techniques to improve the effectiveness and stability of MU methods.
In this work, we first reveal the unstable performance of existing MU methods when they are applied in different scenarios. Considering the need to optimize two objectives on the forgetting and the retaining samples, we then propose Dual Optimizer (DualOptim) to address this challenge. Specifically, we employ an optimizer with adaptive learning rate, such as Adam (kingma2014adam, ; loshchilovdecoupled, ), to optimize the forgetting objective, while utilizing a separate optimizer for the retaining objective. Compared with existing methods, DualOptim decouples the momentum terms for the two objectives by using and updating them separately. In addition, DualOptim is a plug-and-play solution that can be easily integrated into existing MU algorithms.
@@
Figure 1: Unlearning process of MU baselines. SFRon (huang2025unified, ), SalUn (fan2024salun, ) and SCRUB (kurmanji2023towards, ) are adopted as the baselines. The metrics are those mentioned in Sec. 3.1. All results are obtained from unlearning 10% random subset of CIFAR-10 on ResNet-18. The solid lines and shadows denote the mean and standard deviation across 5 trials with different random data. The hyperparameters of different methods are selected based on minimizing the averaging gap between retraining and them across 5 trials. The red dashed lines denote the final performance of retraining as a reference.
Our theoretical analysis validates that decoupling momentum terms in two separate optimizers can help decrease the variance of model parameters, indicating a more stable performance. Through comprehensive experiments, we validate the effectiveness of DualOptim in improving the performance of existing MU methods across various tasks, including image classification, image generation, and large language models. Our proposed technique is generic and capable of pushing the state-of-the-art performance on multiple tasks. We summarize the contributions of this paper as follows:
@@ 2 Related Works
Machine Unlearning (MU). MU targets the need to remove specific data influences from pretrained models (bourtoule2021machine, ; shaik2024exploring, ; xu2024machine, ), while complying with privacy requirements tied to differential privacy (guo2020certified, ; neel2021descent, ; sekhari2021remember, ; ginart2019making, ; ullah2021machine, ). Initially developed on linear classifiers with convex loss objective functions, exact MU (guo2020certified, ; neel2021descent, ; koh2017understanding, ; giordano2019swiss, ) allowed precise data removal under privacy budgets to counter privacy attacks. However, exact MU cannot be applied to deep neural networks due to their non-convex loss functions and prohibitive computational cost for full retraining. In this context, approximate MU (izzo2021approximate, ; thudi2022unrolling, ) was proposed to address this issue by fine-tuning the model to achieve the forgetting effect. Despite improved efficiency, approximate MU may cause catastrophic performance declines on retaining data (thudi2022unrolling, ; golatkar2020eternal, ). Recent methods incorporate fine-tuning (fan2024salun, ; kurmanji2023towards, ; tarun2023fast, ), sparsity regularization (jia2023model, ), knowledge distillation (kurmanji2023towards, ; chundawat2023can, ), saliency map (fan2024salun, ; huang2025unified, ) and alternative updating (huang2025unified, ) to better balance efficient forgetting and model utility. However, as illustrated in Figure 1, these methods generally exhibit suboptimal performance or high hyperparameter sensitivity, underscoring the necessity to develop a method to enhance efficacy and stability during unlearning.
MU for Multiple Tasks. Besides classification, MU has broad applications for multiple tasks. For example, recent text-conditional image generation models have showcased their ability to produce images that closely align with textual descriptions (ho2020denoising, ; ho2021classifier, ; rombach2022high, ; peebles2023scalable, ). However, significant security and privacy concerns have been raised (randored, ; schramowski2023safe, ; carlini2023extracting, ), necessitating the application of MU methods to these models. While early works (schramowski2023safe, ; gandikota2023erasing, ; kumari2023ablating, ; zhang2024forget, ) focuses on concept deletion in diffusion models, recent studies (fan2024salun, ; huang2025unified, ; heng2023selective, ) improve their performance and propose methods applicable to more general image generators. Another notable example is large language models (LLMs), which have demonstrated remarkable capabilities (brown2020language, ; touvron2023llama, ) but also face privacy and copyright issues like retaining unauthorized content (huang2022large, ; carlini2022quantifying, ; staabbeyond, ; dou2024avoiding, ). To address these concerns, MU methods have been employed to fine-tune LLMs (mainitofu, ; yuan2024closer, ; liu2025rethinking, ; zhangnegative, ) to effectively and efficiently achieve data forgetting. However, all these methods struggle to achieve a balance between model utility and forget effectiveness. In addition, intensive hyperparameter tuning is expected for optimal performance, bringing challenges for practitioners.
@@ 3.1 Preliminary
Let $\mathcal{D}=\{{\bm{z}}_{i}\}_{i=1}^{N}$ denote the training set for pretraining. The subset of the training set we aim to forget during unlearning is known as the forget set $\mathcal{D}_{f}\subset\mathcal{D}$, and its complement, $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$ is the retain set. In the context of MU, we denote the parameters of pretrained model as $\theta_{o}\in{\mathbb{R}}^{d}$, which is trained on $\mathcal{D}$. Consistent with previous studies (fan2024salun, ; huang2025unified, ; thudi2022unrolling, ; jia2023model, ), Retraining is considered the gold standard for MU, where the model parameters $\theta_{r}\in\mathbb{R}^{d}$ are trained from scratch on $\mathcal{D}_{r}$. However, retraining is computationally intensive or even infeasible, especially for large models and datasets. This poses a significant challenge in the practical applications of MU.
@@
We focus on approximate MU in this work, its primary objective is to obtain an unlearned model, referred to as $\theta_{u}\in{\mathbb{R}}^{d}$, from $\theta_{o}$ on $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$ so that it can serve as an accurate and computationally efficient alternative to the retrained model $\theta_{r}$. Mathematically, MU aims to solve the optimization problem: $\min_{\theta}\mathcal{L}_{f}(\mathcal{D}_{f},\theta)+\mathcal{L}_{r}(\mathcal{D}_{r},\theta)$, where $\mathcal{L}_{f}$ and $\mathcal{L}_{r}$ denote forget and retain loss objective functions, respectively. In practice, ${\mathcal{L}}_{r}$ is usually the same as the loss objective function in pre-training and ${\mathcal{L}}_{f}$ is the opposite, so MU aims to improve the performance on ${\mathcal{D}}_{r}$ while degrading the performance on ${\mathcal{D}}_{f}$. To avoid confusion, we call ${\mathcal{L}}_{f}({\mathcal{D}}_{f},\theta)$, ${\mathcal{L}}_{r}({\mathcal{D}}_{r},\theta)$ that we aim to minimize forget loss and retain loss, respectively. In addition, many MU algorithms (fan2024salun, ; huang2025unified, ; kurmanji2023towards, ) update $\theta$ by alternately optimizing $\mathcal{L}_{f}({\mathcal{D}}_{f},\theta)$ and $\mathcal{L}_{r}({\mathcal{D}}_{r},\theta)$ for better performance. The generic pipeline of MU is presented in Algorithm 1. For notation simplicity, we use ${\bm{g}}_{f}:=\nabla_{\theta}{\mathcal{L}}_{f}({\mathcal{D}}_{f},\theta)$ and ${\bm{g}}_{r}:=\nabla_{\theta}{\mathcal{L}}_{r}({\mathcal{D}}_{r},\theta)$ to denote the gradients of the forget loss and the retain loss, respectively. In mini-batch updates, we use $\widehat{{\bm{g}}}_{f}$ and $\widehat{{\bm{g}}}_{r}$ to denote the corresponding stochastic gradients.
@@
The numerical analysis in this section is based on classification tasks. We adopt the same evaluation schemes as huang2025unified . That is, we use the accuracy on the forget set (FA), the retain set (RA), test set (TA) and the success rate of membership inference attack (MIA) (song2021systematic, )  [^1] We follow the implementation in huang2025unified and adopt the entropy of the output probabilities as the metric in MIA. on the forgetting set to indicate model utility and the forgetting effectiveness. A competitive unlearning method should have stable performance and small gaps with retraining in all these four evaluation schemes.
@@ 3.2 Adaptive Learning Rate Enables Stable Forgetting
Despite the effectiveness of existing MU methods, achieving satisfactory unlearning performance often relies on carefully selecting hyperparameters. The optimal hyperparameters can vary significantly across different forget sets, complicating the practical applications of MU algorithms. As shown in Figure 2 (a)-(d), we use the same hyperparameters for SFRon (huang2025unified, ), the best-performed baseline in our evaluation, with the default SGD optimizer for 5 trials of different randomly sampled forget sets. We defer more results of other existing methods to Appendix C.2. Despite being intensively tuned, the unified hyperparameters exhibit quite unstable performance across different forget sets during unlearning. This underscores that a hyperparameter configuration effective for certain forget sets may not be suitable for others. When dealing with a new forget set, even if it is sampled in the same way, the need for precise hyperparameter tuning can pose challenges for practitioners.
@@
Figure 2: Unlearning process with different ablations of the proposed method. All results are obtained from unlearning 10% random subset of CIFAR-10 by SFRon (huang2025unified, ) on ResNet-18. (a)-(d) The metrics are those mentioned in Sec. 3.1. The red dashed lines denote the final performance of retraining as a reference. The solid lines and shadows denote the mean and standard deviation across 5 trials with different random forget sets.
Figure 3 (a) demonstrates the magnitude of the gradients on the forget set and the retain set when using SGD optimizer. We can clearly see that the gradients on the forget set have significantly larger magnitudes with substantial variations compared with the ones on the retain set. In this context, it is crucial to adaptively adjust the learning rate to handle different gradient magnitudes. Therefore, we use optimizers with preconditioners, such as Adam (kingma2014adam, ; loshchilovdecoupled, ) to evade tricky hyperparameter tuning for the learning rate. Specifically, Adam employs adaptive learning rates based on the historic gradient magnitudes, making it suitable to handle large gradient magnitude variations during unlearning.
@@
The observations in Figure 2 (a)-(d) and Figure 3 (b) suggest that Adam provides more stable performance and induces smaller gradient norm across different forget sets. However, despite enhanced stability, the noticeable performance gap with retraining, especially in the metric of membership inference attack (MIA), indicates that Adam only achieves suboptimal unlearning efficacy. We need a mechanism to improve performance while ensuring the stability of the method.
@@ 3.3 Decoupled Momentum for Enhanced Stability in Machine Unlearning
Inspired by the disparities in the magnitudes of $\widehat{{\bm{g}}}_{f}$ and $\widehat{{\bm{g}}}_{r}$, along with their non-positive correlation observed in Figure 3 and 4, we introduce decoupled momentum, which employs two separate momentum terms dedicated to the forget loss ${\mathcal{L}}_{f}$ and the retain loss ${\mathcal{L}}_{r}$, respectively. This approach ensures that the momentum update for the forget loss remains unaffected by the gradients of the retaining loss, and vice versa. By decoupling the momentum terms in this manner, we can more effectively optimize the distinct processes of forgetting and retaining, leading to enhanced stability and performance. As illustrated in Figure 2 (a)-(d) (represented by the green lines) and Figure 3 (c), the combination of adaptive learning rate with decoupled momentum not only stabilizes the unlearning process but also leads to improved unlearning efficacy. Note that as presented in Table 5 and Table 6 of Appendix C.2, although other methods like SalUn fan2024salun and SCRUB kurmanji2023towards perform more stable than SFRon, they only achieve suboptimal unlearning performance. Nevertheless, our method can also improve their performance.
@@
(Correlation Bounds) The correlation between the stochastic gradients from the same function in different time steps is bounded while the correlation between stochastic gradients from different functions is non-positive. That is to say, $\exists\tau\in[0,1]$ such that:
@@
The assumption $\forall t_{1},t_{2}$, $\rho(\widehat{{\bm{g}}}_{f,t_{1}},\widehat{{\bm{g}}}_{r,t_{2}})\leq o(\tau)\simeq 0$ is motivated by the observations in Figure 4 of Appendix C.1 and for sake of notation simplicity. Our analyses can be easily extended to the cases if we use a small constant to bound this correlation.
@@
[…]
@@
The proof is deferred to Appendix A.2. Theorem 3.5 formalizes that decoupled momentum induces smaller worst-case parameter variance compared to shared momentum, demonstrating that decoupled momentum theoretically enhances the stability of unlearning. Additionally, we prove that decoupled momentum can also induce smaller worst-case variance of downstream metrics in Appendix A.3.